Boylston Inc.'s total interest bearing debt is $10,000,000 on which it pays interest at the rate of 5%. Boylston's rate of return on assets is 10%. How much cash flow does leverage add to shareholders equity?

A. $500,000

B. $1,000,000

C. $(500,000) negative cash flow

D. $(1,000,000) negative cash flow.
Answer #1

Increase in return from leverage = 10000000*10% = 1,000,000

Less: Interest = 10000000*5% = 500,000

Net cash flow = $500,000